# SafeWork-R1: Coevolving Safety and Intelligence under the AI-45 Law

### v1

@@ 1.2 Technical Roadmap of SafeLadder
Collectively, SafeLadder presents the first unified framework to endow large models with intrinsic safety-oriented thinking through staged optimization, advancing both the capabilities and safety of LLMs. As shown in Fig. 1, we plot the model’s safety and performance scores throughout the staged optimization process. Both safety and performance improve in tandem, achieving a striking $45^{\circ}$ balance rate. This represents a significant step toward building robust, reliable, and trustworthy general-purpose AI.
@@ 3.2.3 Multiobjective Reward Function
Total $\text{Total Reward}=w_{1}R_{\text{visual-focus}}+w_{2}R_{\text{helpful}}+w_{3}R_{\text{format}}+w_{4}R_{\text{task-aware}}$
@@ 3.4 Deliberative Search RL
[…]
@@
2:Initialize $\theta_{0}\in\Theta$, $\lambda_{0}>0$triangleright $\lambda_{0}=0.01$
@@
4:     $g_{\theta}\leftarrow\nabla_{\theta}R(\theta_{k})+\lambda_{k}\nabla_{\theta}U(\theta_{k})$triangleright RL gradients
5:     $\theta_{k+1}\leftarrow\theta_{k}+\alpha_{k}\,g_{\theta}$triangleright — Dual multiplicative-weights step —
6:     $\lambda_{k+1}\leftarrow\lambda_{k}\exp{(\beta(\eta-U(\theta_{k+1})))}$triangleright $\eta=0.9$
@@ 4.2 Human-in-the-Loop Intervention
Firstly, when users are dissatisfied with the reasoning process in the previous LLM response or identify apparent errors, they have the option to manually edit the corresponding text segments. The process of manually modifying the CoT in the $r$-th round of response can be represented as $\hat{C_{r}}\leftarrow\text{HumanEdition}(C_{r})$. The purpose of manual intervention is to enable LLMs to generate a new response different from the original by recognizing user feedback. However, in practice, directly inputting the edited text as part of the dialogue may cause the model to shift attention away from the original query $Q_{r}$ or overlook the modifications, due to inherent limitations in the model structure. To mitigate this, We propose incorporating a new input query in the current dialogue round, while appropriately condensing or discarding the historical context as needed. The new input for the next turn is then formed by combining the original question with this optimized hint as $Q_{r+1}\leftarrow$ concat$(Q_{r},\hat{C_{r}})$. The resulting response, $C_{r+1}$ and $A_{r+1}$, represents the updated reasoning CoT and answer, achieved through this refined human-in-the-loop intervention. Secondly, human intervention was introduced via edit-distance-based corrections to $C_{r}$, interfering with similar memory and enabling genuine reasoning without KVCache reliance zhan2025gthinker . Most incorrect reasoning arises from missing or incorrect step, which affects all subsequent reasoning. During processing, edits are usually focused on this key step, while the rest are either removed or left unchanged. Given the need to trace user modifications to the CoT content within responses, we adapt the Myers Diff myers1986nd algorithm as a foundational approach for implementing fine-grained text change tracking. The methodological framework for this tracking mechanism is outlined as follows. The parameters remain in the same style as above, give the original CoT text $C_{r}$, and after user edit text $\hat{C_{r}}$. $\mathcal{T}(\cdot)$ is the tokenization operation. The text editing of the $k$-th segment is denoted as $\Delta_{k}=\bigl{\langle}s_{k},\ e_{k},\ o_{k},\ \text{text}_{k}\bigr{\rangle}$, which is represented by starting token index $s_{k}$, ending token index $e_{k}$, corresponding operations $o_{k}$. The general operations include no action, deletion, addition, and modification, which can be represented as $o_{k}\in\{\,\text{equal},\ \text{delete},\ \text{insert},\ \text{replace}\,\}$. After computation, the corresponding edition set will be: $\Delta(C_{r},\hat{C_{r}})=\{\Delta_{k}\}_{k=1}^{n}$, where $n$ denotes the number of segmented content.
@@
2:$\mathcal{T}({C_{r}})\leftarrow\text{Tokenize}(C_{r},\text{mode})$
3:$\mathcal{T}({\hat{C_{r}}})\leftarrow\text{Tokenize}(\hat{C_{r}},\text{mode})$
4:$\mathcal{O}\leftarrow\text{SequenceMatcher}(\mathcal{T}_{A},\mathcal{T}_{B})$
5:$\Delta(C_{r},\hat{C_{r}})\leftarrow\bigl{\{}\langle s_{k},e_{k},o_{k},\text{text}_{k}\rangle\bigm{|}o_{k}\neq\text{equal}\bigr{\}}$
6:return $\Delta(C_{r},\hat{C_{r}}),\ \mathcal{D}(C_{r},\hat{C_{r}})$
@@
5:$N\leftarrow 4$triangleright Set maximum allowed consecutive failures
@@
7:     $Q^{s^{\prime}}\leftarrow$Response$(Q^{s})$triangleright Simplify question by LLM
8:     $(C^{\prime},A^{\prime})\leftarrow$Response by SafeWork-R1$(Q^{s^{\prime}})$triangleright Answer questions by LLM
@@
10:         $Q^{s}\leftarrow Q^{s^{\prime}}$triangleright Update it with new version
[…]
@@
[…]
@@
16:return $Q^{s}$triangleright Return the final simplified question
@@ 5.1 Safety Evaluation
Model MM-SafetyBench MSSBench XSTest-Safe SIUO Avg.
Gemini 2.5 pro     79.3$\uparrow$21.3     70.5$\uparrow$21.0     100.0$\uparrow$8.0     76.7$\uparrow$42.3     81.6$\uparrow$23.2
Claude Opus 4     82.1$\uparrow$21.3     59.6$\uparrow$21.0     96.8$\uparrow$8.0     62.8$\uparrow$42.3     75.3$\uparrow$23.2
GPT-4.1     78.2$\uparrow$21.3     69.1$\uparrow$21.0     96.4$\uparrow$8.0     92.9$\uparrow$42.3     84.1$\uparrow$23.2
GPT-4o     70.2$\uparrow$21.3     58.8$\uparrow$21.0     94.0$\uparrow$8.0     51.8$\uparrow$42.3     68.7$\uparrow$23.2
Qwen2.5-VL-72B     70.4$\uparrow$21.3     53.8$\uparrow$21.0     91.2$\uparrow$8.0     38.2$\uparrow$42.3     63.4$\uparrow$23.2
SafeWork-R1     92.0$\uparrow$21.6     74.8$\uparrow$21.0     99.2$\uparrow$8.0     90.5$\uparrow$52.3     89.2$\uparrow$25.8
@@ 5.2 Value Evaluation
Table 7: Performance of models on value benchs.
@@
Gemini 2.5 Pro     16.8$\uparrow$26.2     70.0$\uparrow$9.7     66.2$\uparrow$6.5     86.8$\downarrow$4.8     44.7$\uparrow$15.0
Claude Opus 4     38.1$\uparrow$26.2     70.7$\uparrow$9.7     74.7$\uparrow$6.5     72.5$\downarrow$4.8     52.2$\uparrow$15.0
GPT-4.1     33.3$\uparrow$26.2     74.4$\uparrow$9.7     62.7$\uparrow$6.5     61.7$\downarrow$4.8     53.0$\uparrow$15.0
GPT-4o     36.6$\uparrow$26.2     72.4$\uparrow$9.7     65.9$\uparrow$6.5     79.7$\downarrow$4.8     55.5$\uparrow$15.0
Qwen2.5-VL-72B     39.1$\uparrow$26.2     58.4$\uparrow$9.7     48.1$\uparrow$6.5     75.7$\downarrow$4.8     49.9$\uparrow$15.0
Safework-R1     65.3$\uparrow$26.2     68.1$\uparrow$9.7     54.6$\uparrow$6.5     70.9$\downarrow$4.8     64.9$\uparrow$15.0
@@ 5.5 Search with Calibration
The comprehensive results are shown in Table 10. While it is challenging to match the latest proprietary SOTA models in accuracy metrics using a base model released six months ago, we maintain substantial advantages in reliability, particularly in the FC(False-Certain ratio) metric.
@@
Table 10: Search with calibration evaluation results. Relib. (↑) abbreviates reliability (consistency between model confidence and correctness). FC(↓) represents the proportion of False yet Certain responses, which is the least desirable scenario for users.
@@
Acc. Relib. FCAcc. Relib. FCAcc. Relib. FCAcc. Relib. FCAcc. Relib. FC%
@@ 5.6 Evaluation and Analysis on General Benchmark
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-MM Avg.
Gemini 2.5 Pro     82.0$\uparrow$3.7     83.0$\uparrow$1.3     81.8$\uparrow$19.5     86.9$\uparrow$9.1     87.2$\uparrow$5.1     84.2$\uparrow$7.7
Claude Opus 4     73.0$\uparrow$3.7     73.0$\uparrow$1.3     68.5$\uparrow$19.5     74.7$\uparrow$9.1     73.7$\uparrow$5.1     72.6$\uparrow$7.7
GPT-4.1     72.4$\uparrow$3.7     72.0$\uparrow$1.3     49.0$\uparrow$19.5     69.2$\uparrow$9.1     60.2$\uparrow$5.1     64.6$\uparrow$7.7
GPT-4o     70.6$\uparrow$3.7     61.6$\uparrow$1.3     33.7$\uparrow$19.5     46.9$\uparrow$9.1     33.8$\uparrow$5.1     49.3$\uparrow$7.7
Qwen2.5-VL-72B     67.2$\uparrow$3.7     74.8$\uparrow$1.3     40.4$\uparrow$19.5     50.5$\uparrow$9.1     73.1$\uparrow$5.1     61.2$\uparrow$7.7
Safework-R1     70.9$\uparrow$3.7     76.1$\uparrow$1.3     59.9$\uparrow$19.5     59.6$\uparrow$9.1     78.2$\uparrow$5.1     68.9$\uparrow$7.7
@@ 7 Conclusions and Discussions
Chao Yang Xingcheng Xu Yizhuo Ding
@@
Bo Zhang Chen Qian Dongxing Shi Fenghua Weng Guanxu Chen
Hao Li Hefeng Zhou Hong Huang Jiashu Qu Jinxuan Zhang
Juncheng Li Lingxiao Du Lujun Gui Mingkang Chen Pengyu Zhu
Qianxi He Qihan Ren Qihua Liu Shuai Shao Shujie Wang
Tianle Gu Wanying Qu Xiangtian Li Xin Song Xin Wang
Xinhao Song Xinshun Feng Yafu Li Yang Yao Yige Wang
Yixu Wang Yuan Pu Yuanfu Wang Yudong Lu Yue Yang
Yuenan Hou Yulei Ye Yunhao Chen Yuqi Wu Zhenyun Yin
Zhixuan Liu Ziqi Miao
@@ Appendix A Appendix: Evaluation on Various Models
Model MM-SafetyBench MSSBench XSTest-Safe SIUO FLAMES M3oralBench
Qwen2.5-VL-7B     50.1$\uparrow$38.2     51.7$\uparrow$23.4     96.8$\uparrow$2.0     30.8$\uparrow$53.4     32.4$\uparrow$32.7     51.1$\uparrow$9.4
SafeWork-R1-Qwen2.5VL-7B     88.3$\uparrow$38.2     65.1$\uparrow$23.4     98.8$\uparrow$2.0     84.2$\uparrow$53.4     65.1$\uparrow$32.7     60.5$\uparrow$9.4
@@
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-MM
Qwen2.5-VL-7B     49.6$\uparrow$6.3     66.2$\uparrow$5.0     23.2$\uparrow$4.3     30.3$\uparrow$0.0     51.2$\uparrow$15.0
SafeWork-R1-Qwen2.5VL-7B     55.9$\uparrow$6.3     71.2$\uparrow$5.0     27.5$\uparrow$4.3     30.3$\uparrow$0.0     76.2$\uparrow$25.0
@@ A.2 Experiment on InternVL3-78B
Model MM-SafetyBench MSSBench XSTest-Safe SIUO FLAMES M3oralBench
InternVL3-78B     71.0$\uparrow$17.6     52.8$\uparrow$22.6     100.0$\downarrow$1.2     44.2$\uparrow$42.1     32.3$\uparrow$25.6     68.2$\uparrow$3.9
SafeWork-R1-InternVL3-78B     88.6$\uparrow$17.6     75.4$\uparrow$22.6     98.8$\downarrow$1.2     86.3$\uparrow$42.1     57.8$\uparrow$25.6     72.0$\uparrow$3.9
@@
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-MM
InternVL3-78B     67.3$\uparrow$0.3     74.3$\uparrow$0.1     44.6$\uparrow$8.2     48.5$\uparrow$8.6     69.7$\uparrow$2.2
SafeWork-R1-InternVL3-78B     67.7$\uparrow$0.4     74.4$\uparrow$0.1     52.8$\uparrow$8.2     57.1$\uparrow$8.6     71.8$\uparrow$2.1
@@ A.3 Experiment on DeepSeek-R1-Distill-Llama-70B
DeepSeek-R1-Distill-Llama-70B     96.8$\uparrow$1.2     21.8$\downarrow$21.3     62.0$\downarrow$61.8     31.6$\uparrow$40.6     69.5$\uparrow$29.8
SafeWork-R1-DeepSeek-70B     98.0$\uparrow$1.2     0.5$\downarrow$21.3     0.2$\downarrow$61.8     72.2$\uparrow$40.6     99.3$\uparrow$29.8
@@
DeepSeek-R1-Distill-Llama-70B     59.1$\downarrow$1.0     93.2$\downarrow$1.4     67.1$\uparrow$7.1     41.9$\uparrow$8.6     40.0$\uparrow$8.0
SafeWork-R1-DeepSeek-70B     58.1$\downarrow$1.0     91.8$\downarrow$1.4     74.2$\uparrow$7.1     50.5$\uparrow$8.6     48.0$\uparrow$8.0

### v3

@@ 1.2 Technical Roadmap of SafeLadder
Collectively, SafeLadder presents the first unified framework to endow large models with intrinsic safety-oriented thinking through staged optimization, advancing both the capabilities and safety of LLMs. As shown in Fig. 1, we plot the model’s safety and performance scores throughout the staged optimization process. Both safety and performance improve in tandem, achieving the AI-$45^{\circ}$ Law. This represents a significant step toward building robust, reliable, and trustworthy general-purpose AI.
@@ 3.2.3 Multiobjective Reward Function
$\text{Total Reward}=w_{1}R_{\text{visual-focus}}+w_{2}R_{\text{helpful}}+w_{3}R_{\text{format}}+w_{4}R_{\text{task-aware}}$
@@ 3.4 Deliberative Search RL
$P^{*}=\underset{\theta}{\max}\ \underset{\lambda\geq 0}{\min}\ \mathcal{L}(\theta,\lambda)=R(\theta)+\sum_{i=1}^{m}\lambda_{i}(U_{i}(\theta)-\eta_{i}),$ (6)
Since paternain2022safe demonstrated the strong duality holds for Eq.˜6 under the setting of RL, we only need to solve: $Q^{*}=\underset{\lambda\geq 0}{\min}\ \underset{\theta}{\max}\ \mathcal{L}(\theta,\lambda)$. Our dynamic RL algorithm can be formulated as follows:
@@
2:Initialize $\theta_{0}\in\Theta$, $\lambda_{0}>0$ $\triangleright$ $\lambda_{0}=0.01$
@@
4:  $g_{\theta}\leftarrow\nabla_{\theta}R(\theta_{k})+\lambda_{k}\nabla_{\theta}U(\theta_{k})$ $\triangleright$ RL gradients
5:  $\theta_{k+1}\leftarrow\theta_{k}+\alpha_{k}\,g_{\theta}$ $\triangleright$ — Dual multiplicative-weights step —
6:  $\lambda_{k+1}\leftarrow\lambda_{k}\exp{(\beta(\eta-U(\theta_{k+1})))}$ $\triangleright$ $\eta=0.9$
@@ 4.2 Human-in-the-Loop Intervention
[…]
@@
[…]
@@
[…]
6:while $\text{fail\_count}<N$ do
7:  $Q^{s^{\prime}}\leftarrowResponse$$(Q^{s})$ $\triangleright$ Simplify question by LLM
[…]
@@
10:   $Q^{s}\leftarrow Q^{s^{\prime}}$ $\triangleright$ Update it with new version
11:   $\text{fail\_count}\leftarrow 0$ $\triangleright$ Reset fail counter
@@
13:   $\text{fail\_count}\leftarrow\text{fail\_count}+1$ $\triangleright$ Increase fail counter
@@
16:return $Q^{s}$ $\triangleright$ Return the final simplified question
@@ 5.1 Safety Evaluation
Model $M$-SafetyBench MSSBench XSTest-Safe SIUO Avg.
[…]
@@ 5.2 Value Evaluation
Table 7: Performance of models on value benchmarks.
@@
[…]
@@ 5.5 Search with Calibration
The comprehensive results are shown in Table 10. While it is challenging to match the latest proprietary SOTA models in accuracy metrics using a base model released six months ago, we maintain substantial advantages in reliability, particularly in the FC% (False-Certain ratio) metric.
@@
Table 10: Search with calibration evaluation results. Relib. (↑) abbreviates reliability (consistency between model confidence and correctness). FC% (↓) represents the proportion of False yet Certain responses, which is the least desirable scenario for users.
@@
Acc. Relib. FC% Acc. Relib. FC% Acc. Relib. FC% Acc. Relib. FC% Acc. Relib. FC%
@@ 5.6 Evaluation and Analysis on General Benchmark
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-$M$ Avg.
[…]
@@ 7 Conclusions and Discussions
Bo Zhang Chao Yang Xingcheng Xu Yizhuo Ding
@@
Chen Qian Dongxing Shi Fenghua Weng Guanxu Chen Hao Li
Hefeng Zhou Hong Huang Jiashu Qu Jinxuan Zhang Juncheng Li
Lingxiao Du Lujun Gui Mingkang Chen Pengyu Zhu Qianxi He
Qihan Ren Qihua Liu Shuai Shao Shujie Wang Tianle Gu
Wanying Qu Xiangtian Li Xin Song Xin Wang Xinhao Song
Xinshun Feng Yafu Li Yang Yao Yige Wang Yixu Wang
Yuan Pu Yuanfu Wang Yudong Lu Yue Yang Yuenan Hou
Yulei Ye Yunhao Chen Yuqi Wu Zhenyun Yin Zhixuan Liu
Ziqi Miao
@@ Appendix A Appendix: Evaluation on Various Models
Model $M$-SafetyBench MSSBench XSTest-Safe SIUO FLAMES M3oralBench
[…]
@@
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-$M$
[…]
@@ A.2 Experiment on InternVL3-78B
Model $M$-SafetyBench MSSBench XSTest-Safe SIUO FLAMES M3oralBench
[…]
@@
Model MMMU MathVista Olympiad GPQA Diamond GAOKAO-$M$
[…]
@@ A.3 Experiment on DeepSeek-R1-Distill-Llama-70B
[…]
@@
[…]